OpenStudy (anonymous):

solve for x 6x^2+13x=63

OpenStudy (anonymous):

(2x - 9)(3x + 7)

OpenStudy (anonymous):

needed in fraction form

OpenStudy (anonymous):

ok

OpenStudy (shamim):

\[6x ^{2}+13x=63 \]

OpenStudy (shamim):

\[6x ^{2}+13x-63=0\]

OpenStudy (shamim):

\[6x ^{2}+27x-14x-63=0\]

OpenStudy (shamim):

r u following me

OpenStudy (anonymous):

yes i understand

OpenStudy (shamim):

\[or, 3x(2x+9)-7(2x+9)=0\]

OpenStudy (shamim):

so i can stop now

OpenStudy (shamim):

is not it

OpenStudy (highlyeducated87):

The answer in fraction is x=-9/2 , 1/3

OpenStudy (shamim):

no

OpenStudy (shamim):

please follow me

OpenStudy (shamim):

\[or, (3x-7)(2x+9)=0\]

OpenStudy (shamim):

\[or, 3x-7=0\]

OpenStudy (shamim):

\[or, x=\frac{ 7 }{ 3 }\]

OpenStudy (shamim):

another answer will be as bellow

OpenStudy (shamim):

\[2x+9=0\]

OpenStudy (shamim):

\[or, x=-\frac{ 9 }{ 2 }\]

OpenStudy (shamim):

i think u got my solution